# NuGraph2: A Graph Neural Network for Neutrino Physics Event Reconstruction

### v1

@@ II Experimental setup
MicroBooNE is a LArTPC detector operating at Fermilab [20], consisting of a cryostat containing cryogenically cooled liquid argon. Charged particles interacting in the argon volume produce ionization electrons that drift under an electric field to a series of three wire planes; these electrons induce electronic signals on the first two wire planes before being collected on the third. The three wire planes (U, V and Y) are oriented at an angle of $+60\deg$, $-60\deg$ and $0\deg$ with respect to vertical, respectively. If the $t_{0}$ of a particle interaction is known, then the time at which signal pulses on TPC wires occur can be used as a measure of the signal’s distance from the wires. Due to the differing wire angles, each TPC wire plane provides a unique 2D representation of a particle interaction; these three 2D representations can be combined to provide a full 3D reconstruction of a particle interaction.
In order to train the NuGraph2 architecture, this dataset is used to construct heterogeneous graph objects representing each neutrino interaction; for this paper, only hits passing the Pandora [21] neutrino slice selection are considered. Pulses on TPC wires are summarized into discrete Gaussian hits occurring on specific wires at specific times. An independent subgraph is formed for each wire plane, with the hits on a given plane used as the basis for that plane’s graph nodes. Four input features are constructed for each graph node: the hit’s wire index, time coordinate, and the integral and Root Mean Square (RMS) width of its Gaussian pulse.
@@ III.1 Encoder
NuGraph2’s encoder serves to take the input graph node features $x^{\textrm{in}}$ and generate the initial categorical node embedding. First, the $(N_{n},N_{f})$ input tensor is simply repeated $N_{c}$ times across a new second dimension to produce a new tensor $x^{c}$ with shape $(N_{n},N_{c},N_{f})$. This is then used to generate the initial node embedding
@@
where $\phi_{e}$ is a categorical linear multi-layer perceptron (MLP).
@@ III.2 Message passing
At the start of each message-passing iteration, graph node features at iteration $l$, $n^{l}_{i}$, are concatenated with $x^{c}$, the input graph node features prior to any convolutions repeated across the categorical dimension, to form the $x_{i}^{l}$ tensor.
@@ III.2.1 Planar block
where $\phi_{p}$ is a categorical linear MLP and $p^{l}_{i}$ are updated planar node features.
@@ III.2.2 Nexus block
where $k=\mathcal{X}(i)$ are the nexus nodes connected to planar node $i$, and $\theta_{u}$ is a simple message-passing convolution operator that transforms planar node features to nexus nodes. By construction each nexus node is connected to a maximum of one node on each plane, and so planar features are aggregated into nexus nodes via concatenation. In the event that a nexus node is not connected to any node on a given plane, that plane’s contribution to the nexus node’s features are set to zero.
Nexus node features are then convolved to produce new nexus features
@@ IV Training
NuGraph2 utilizes PyTorch [27] with PyG [28] for our model, leveraging the model training infrastructure provided by PyTorch Lightning. Training is optimized using the AdamW optimizer [29] with a learning rate of 0.001, scheduled using the OneCycleLR scheduler [30] for a total of 80 epochs.
@@ V.1 Model performance
For the network’s semantic probability outputs, a discrete class prediction per hit is determined by selecting the class with the largest probability score. The semantic decoder achieves an overall recall and precision of 0.949; these values are equal due to the scheme used to weight the loss function. Recall and precision confusion matrices are shown in Fig. 4.
@@
The shower, Michel and diffuse classes have similar conceptual overlap, as all three more frequently consist of smaller energy depositions – electromagnetic showers are typically formed from a cascade of smaller and frequently disconnected energy depositions, while Michel electrons are short tracks, and diffuse activity often manifests as small and disconnected energy depositions due to activity from neutrons and EM activity such as Compton scatters. This is reflected in the network output, as 11% and 10% of true Michel hits are misclassified as shower-like and diffuse respectively. Similarly, 5.8% and 2.1% of true diffuse hits are misclassified as shower-like and Michel-like, respectively.
@@ VI Summary
The NuGraph2 architecture is a state-of-the-art graph neural network for particle reconstruction in neutrino physics, achieving 98.0% efficiency in filtering background hits and 94.9% efficiency in semantically labelling hits, while producing 2D representations that are 94.8% consistent. Unlike CNN-based techniques, it is capable of operating on the detector observables in their native structure, without requiring any arbitrary pixelization, voxelization, truncation or downsizing.

### v2

# Graph Neural Network for Neutrino Physics Event Reconstruction

@@ II Experimental setup
MicroBooNE is a LArTPC detector operating at Fermilab [20], consisting of a cryostat containing cryogenically cooled liquid argon. Charged particles interacting in the argon volume produce ionization electrons that drift under an electric field to a series of three wire planes; these electrons induce electronic signals on the first two wire planes before being collected on the third. The three wire planes (U, V and Y) are oriented at an angle of $+60^{\circ}$, $-60^{\circ}$ and $0^{\circ}$ with respect to vertical, respectively. If the $t_{0}$ of a particle interaction is known, then the time at which signal pulses on TPC wires occur can be used as a measure of the signal’s distance from the wires. Due to the differing wire angles, each TPC wire plane provides a unique 2D representation of a particle interaction; these three 2D representations can be combined to provide a full 3D reconstruction of a particle interaction.
In order to train the NuGraph2 architecture, this dataset is used to construct heterogeneous graph objects representing each neutrino interaction. For this paper, only hits passing the Pandora [21] neutrino slice selection are considered. This selection is based on 3D reconstruction and proximity and direction-based metrics, and separates 2D hits into slices representing distinct neutrino or cosmic ray interactions. Pulses on TPC wires are summarized into discrete Gaussian hits occurring on specific wires at specific times. An independent subgraph is formed for each wire plane, with the hits on a given plane used as the basis for that plane’s graph nodes. Four input features are constructed for each graph node: the hit’s wire index, time coordinate, and the integral and Root Mean Square (RMS) width of its Gaussian pulse.
@@
Planar graph nodes from the simulated neutrino interaction are assigned true semantic labels based on underlying simulation truth. The semantic labelling scheme considers five classes: MIP, which refers to particle tracks from minimum ionizing particles such as muons and charged pions; HIP, which refers to particle tracks from highly ionizing particles such as protons and charged kaons; shower, which refers to cone-shaped electromagnetic cascades produced by photons or electrons; Michel, which refers to Michel electrons produced at the end of a muon track; and diffuse, which refers to any particles with low hit multiplicity that do not produce an identifiable object, such as neutrons or electromagnetic activity due to Compton scattering.
@@ III.1 Encoder
NuGraph2’s encoder serves to take the input graph node features $x^{\textrm{in}}$ and generate the initial categorical node embedding. First, the $(N_{n},N_{i})$ input tensor is repeated $N_{c}$ times across a new second dimension to produce a new feature embedding on the planar nodes $x^{c}$ with shape $(N_{n},~{}N_{c},~{}N_{i})$, where $N_{i}$ is the number of node features. This is then used to generate the initial node embedding
@@
[…]
@@ III.2 Message passing
[…]
@@ III.2.1 Planar block
[…]
@@ III.2.2 Nexus block
where $k=\mathcal{X}(i)$ are the nexus nodes connected to planar node $i$, and $\theta_{u}$ is a simple message-passing convolution operator that transforms planar node features to nexus nodes. Since nexus nodes are derived from spacepoints that group together hits across detector planes, each nexus node is connected to a maximum of one planar node per plane, and so planar features are aggregated into nexus nodes via concatenation. In the event that a nexus node is not connected to any node on a given plane, that plane’s contribution to the nexus node’s features are set to zero.
Features on each nexus node are then passed forward into a categorical MLP to produce new nexus features
@@ IV Training
The graph dataset consists of 333771 events, which are subdivided according to a 90%/5%/5% split into training, validation and testing samples containing 300395, 16688 and 16688 events respectively. At the start of each epoch, the training dataset is shuffled at random into 4693 batches of 64 events, with each of these batches forming the input to a single iteration of training.
NuGraph2 utilizes PyTorch [27] with PyG [28] for our model, leveraging the model training infrastructure provided by PyTorch Lightning. Training is optimized using the AdamW optimizer [29] with a learning rate of 0.001, scheduled using the OneCycleLR scheduler [30] for a total of 80 epochs (375,440 iterations).
@@ V.1 Model performance
For the network’s semantic probability outputs (using the labelling scheme defined in Sec. II), a discrete class prediction per hit is determined by selecting the class with the largest probability score. The semantic decoder achieves an overall recall and precision of 0.949; these values are equal due to the scheme used to weight the loss function. Recall and precision confusion matrices are shown in Fig. 4.
@@
The shower, Michel and diffuse classes have similar conceptual overlap, as all three more frequently consist of smaller energy depositions – electromagnetic showers are typically formed from a cascade of smaller and frequently disconnected energy depositions, while Michel electrons are short tracks, and diffuse activity often manifests as small and disconnected energy depositions. This is reflected in the network output, as 11% and 10% of true Michel hits are misclassified as shower-like and diffuse respectively. Similarly, 5.8% and 2.1% of true diffuse hits are misclassified as shower-like and Michel-like, respectively.
@@ VI Summary
NuGraph2 is a graph neural network architecture for particle reconstruction in neutrino physics, achieving 98.0% efficiency in filtering background hits and 94.9% efficiency in semantically labelling hits, while producing 2D representations that are 94.8% consistent. Unlike CNN-based techniques, it is capable of operating on the detector observables in their native structure, without requiring any arbitrary pixelization, voxelization, truncation or downsizing.